### v1

@@ 5 Area-Weighted-Peeling Algorithm
To restore the broken part of $L^{2}$, we perform extreme point queries on the remaining points efficiently using a dynamic convex hull data structure $D_{CH}$ (line 1) as in [18] or [6]. As described in Lemma 6.9, $A(u)$ is always contiguous on $L^{2}$. Therefore, removing $A(u)$ from $L^{2}$ requires us to restore it between two ”endpoints” $a$ and $b$. The first extreme point query uses line $\overline{ab}$ in the direction of $u$. If a point $z$ from $D_{CH}$ is found then at least two more queries are performed with lines $\overline{za}$ and $\overline{zb}$. In general, if $k$ points are found then the number of queries is $12k+$ 1. The $k$ points are deleted from $D_{CH}$. This all happens on line 1.
@@ 5.1 Analysis
The hardest part of the analysis is showing that the overall time spent on lines 1 and 1 is $\mathcal{O}\!\left(n\log n\right)$. We first show that, excluding the time spent on these lines, the running time of Algorithm 1 is $\mathcal{O}\!\left(n\log{n}\right)$. In line 1 we compute the first and second convex layers in $\mathcal{O}\!\left(n\log{n}\right)$ time by running any optimal convex hull algorithm twice. In lines 1 to 1, we compute the initial sensitivities by finding the points active for each $u\in L^{1}$. As described above, we can do this by applying two tangent queries, allowing us to recover the first and last extreme point for $u$. We can walk along $L^{2}$ between them to recover $A(u)$. Once $A(u)$ is found for each $u$, we find $\sigma\!\left(u\right)$ by computing the area of the polygon $\pentagon\!\left(t\circ u\circ v\circ A(u)\right)$, where $t$ and $v$ are $u$’s neighbors. By Lemma 6.1(1), in this initial configuration each point on $L^{2}$ is active in at most three triangles. Thus, we make in total $\mathcal{O}\!\left(|L^{2}|\right)=\mathcal{O}\!\left(n\right)$ tangent queries, each of which costs $\mathcal{O}\!\left(\log n\right)$ time. Since the area of a simple polygon can be computed in linear time [26], all the area computations take $\sum_{u\in L^{1}}\Theta\!\left(1+\left|A(u)\right|\right)=\mathcal{O}\!\left(\left|L^{1}\right|+\left|L^{2}\right|\right)=\mathcal{O}\!\left(n\right)$ time. Therefore, the overall time to initialize the priority queue is $\mathcal{O}\!\left(n\log n\right)$.
@@ 6.1 Preliminaries
    Let $p$ be a point on any layer $k$. After deleting any point $q\neq p$ and reconstructing the convex layers, $p$ is on layer $k-1$ or $k$.
@@ 6.2 Bounding the active points
First, we consider the case where $v^{\prime}\notin\triangle tuv$. We want to show that $p\in\triangle tuv^{\prime}$. Equivalently, that $p$ is in the intersection of the three half-planes $\overrightarrow{tu}$, $\overrightarrow{uv^{\prime}}$, and $\overrightarrow{tv^{\prime}}$. Clearly, $p$ must satisfy the half-planes $\overrightarrow{tu}$ and $\overrightarrow{uv^{\prime}}$ as these coincide with hull edges. In addition, since $v^{\prime}\notin\triangle tuv$, the half-plane for $\overrightarrow{tv}$ is a subset of the half-plane for $\overrightarrow{tv^{\prime}}$. Because $p\in\triangle tuv$, $p$ satisfies $\overrightarrow{tv}$. Therefore, $p$ must satisfy $\overrightarrow{tv^{\prime}}$.
Now we consider the case where $v^{\prime}\in\triangle tuv$. Assume that $p\notin\triangle tuv^{\prime}$. Then because we know that $p\in\triangle tuv$, either $p\in\triangle tv^{\prime}v$ or $p\in\triangle uv^{\prime}v$. If $p\in\triangle tv^{\prime}v$, by Lemma 6.1(2), $p$ could not have been active for $u$ prior to deleting $v$. If $p\in\triangle uv^{\prime}v$, $p$ is now outside of the convex hull. Either way, this is a contradiction.
@@ 6.3 Updating sensitivities
Let $(u,v)$ be points on the first layer. Consider a peel of $v$ where $\delta_{u}$ new points become active points for $u$. Then the updated sensitivity $\sigma\!\left(u\right)$ can be computed it $\Theta\!\left(\delta_{u}+\log n\right)$ time, excluding the time to restore the second and first layer.
@@ Acknowledgement
Let $p$ be a point on any convex layer $k$. After deleting any point $q\neq p$ and reconstructing the convex layers, $p$ is on layer $k-1$ or $k$.
@@
Now we will show that $p$ never moves up more than one layer at a time. This is clearly true for $L^{1}$ and $L^{2}$ because only one point is completely removed from the structure at at time (i.e. shifts to layer 0). For layers $k\geq 3$, consider a point $p$ on layer $k$ that moves to layer $k^{\prime}\leq k-2$. Let $L^{*k^{\prime}}$ be the set of points on layer $k^{\prime}$ after deleting $q$. Let $L^{k-1}$ be the set of points on layer $k-1$ before deleting $q$.
Because $p\in L^{*k^{\prime}}$, no convex combination of the points in $L^{*k^{\prime}}-\{p\}$ equals $p$ by convexity. By the inductive hypothesis, all points on $L^{k-1}$ are convex combinations of $L^{*k^{\prime}}$ because upon deleting $q$ no point on $L^{k-1}$ advances above layer $k^{\prime}$. Furthermore, they are all convex combinations of $L^{*k^{\prime}}-\{p\}$ as $p$ itself is a convex combination of $L^{k-1}$. But if $p$ is not a convex combination of $L^{*k^{\prime}}-\{p\}$, and all the points on layer $k-1$ are convex combinations of $L^{*k^{\prime}}-\{p\}$, then prior to deleting $q$, $p$ was above layer $k-1$, which is a contradiction.

### v2

@@ 5 Area-Weighted-Peeling Algorithm
To restore the broken part of $L^{2}$, we perform extreme point queries on the remaining points efficiently using a dynamic convex hull data structure $D_{CH}$ (line 1) as in [18] or [6]. As described in Lemma 6.9, $A(u)$ is always contiguous on $L^{2}$. Therefore, removing $A(u)$ from $L^{2}$ requires us to restore it between two “endpoints” $a$ and $b$. The first extreme point query uses line $\overline{ab}$ in the direction of $u$. If a point $z$ from $D_{CH}$ is found then at least two more queries are performed with lines $\overline{za}$ and $\overline{zb}$. In general, if $k$ points are found then the number of queries is $2k+1$. The $k$ points are deleted from $D_{CH}$. This all happens on line 1.
@@ 5.1 Analysis
The hardest part of the analysis is showing that the overall time spent on lines 1 and 1 is $\mathcal{O}\!\left(n\log n\right)$. We first show that, excluding the time spent on these lines, the running time of Algorithm 1 is $\mathcal{O}\!\left(n\log{n}\right)$. In line 1 we compute the first and second convex layers in $\mathcal{O}\!\left(n\log{n}\right)$ time by running any optimal convex hull algorithm twice. In lines 1 to 1, we compute the initial sensitivities by finding the points active for each $u\in L^{1}$. As described above, we can do this using two tangent queries on $L^{2}$ from the neighbors of $u$. Once $A(u)$ is found for each $u$, we find $\sigma\!\left(u\right)$ by computing the area of the polygon $\pentagon\!\left(t\circ u\circ v\circ A(u)\right)$, where $t$ and $v$ are the neighbors of $u$. In total, we make $\mathcal{O}\!\left(|L^{1}|\right)=\mathcal{O}\!\left(n\right)$ tangent queries, each of which takes $\mathcal{O}\!\left(\log n\right)$ time. Since the area of a simple polygon can be computed in linear time [26], and by Lemma 6.1(1) each point on $L^{2}$ is active in at most three triangles in the initial configuration, the total time to compute all areas is $\sum_{u\in L^{1}}\Theta\!\left(1+\left|A(u)\right|\right)=\mathcal{O}\!\left(\left|L^{1}\right|+\left|L^{2}\right|\right)=\mathcal{O}\!\left(n\right)$ time. Therefore, the overall time to initialize the priority queue is $\mathcal{O}\!\left(n\log n\right)$.
@@ 6.1 Preliminaries
    Let $p$ be a point on any layer $k$. After deleting any point $q\neq p$ and reconstructing the convex layers, $p$ is on layer k1k-1 or $k$.
@@ 6.2 Bounding the active points
First, we consider the case where $v^{\prime}\notin\triangle tuv$. We want to show that $p\in\triangle tuv^{\prime}$. Equivalently, that $p$ is in the intersection of the three half-planes $\overrightarrow{tu}$, $\overrightarrow{uv^{\prime}}$, and $\overrightarrow{tv^{\prime}}$. Clearly, $p$ must satisfy the half-planes $\overrightarrow{tu}$ and $\overrightarrow{uv^{\prime}}$ as these coincide with hull edges. In addition, since $v^{\prime}\notin\triangle tuv$, the half-plane for $\overrightarrow{tv}$ is a subset of the half-plane for $\overrightarrow{tv^{\prime}}$. Because $\in\triangle tuv, pp$ satisfies $\overrightarrow{tv}$. Therefore, $p$ must satisfy $\overrightarrow{tv^{\prime}}$.
Now we consider the case where $v^{\prime}\in\triangle tuv$. Assume that $p\notin\triangle tuv^{\prime}$. Then because we know that $p\in\triangle tuv$, either $p\in\triangle tv^{\prime}v$ or $p\in\triangle uv^{\prime}v$. If $p\in\triangle tv^{\prime}v$, by Lemma 6.1(2), $p$ could not have been active for $u$ prior to deleting $v$. If $\in\triangle uv^{\prime}v, pp$ is now outside of the convex hull. Either way, this is a contradiction.
@@ 6.3 Updating sensitivities
Let $(u,v)$ be points on the first layer. Consider a peel of $v$ where $\delta_{u}$ new points become active points for $u$. Then the updated sensitivity $\sigma\!\left(u\right)$ can be computed in $\Theta\!\left(\delta_{u}+\log n\right)$ time, excluding the time to restore the second and first layer.
@@ Acknowledgement
Let $p$ be a point on any convex layer $k$. After deleting any point $q\neq p$ and reconstructing the convex layers, $p$ is on layer k1k-1 or $k$.
@@
Now we will show that $p$ never moves up more than one layer at a time. This is clearly true for $L^{1}$ and $L^{2}$ because only one point is completely removed from the structure at at time (i.e. shifts to layer 0). For layers $k\geq 3$, consider a point $p$ on layer $k$ that moves to layer $k^{\prime}\leq k-2$. Let $L^{*k^{\prime}}$ be the set of points on layer $k^{\prime}$ after deleting $q$. Let $L^{k-1}$ be the set of points on layer k1k-1 before deleting $q$.
Because $p\in L^{*k^{\prime}}$, no convex combination of the points in $L^{*k^{\prime}}-\{p\}$ equals $p$ by convexity. By the inductive hypothesis, all points on $L^{k-1}$ are convex combinations of $L^{*k^{\prime}}$ because upon deleting $q$ no point on $L^{k-1}$ advances above layer $k^{\prime}$. Furthermore, they are all convex combinations of $L^{*k^{\prime}}-\{p\}$ as $p$ itself is a convex combination of $L^{k-1}$. But if $p$ is not a convex combination of $L^{*k^{\prime}}-\{p\}$, and all the points on layer k1k-1 are convex combinations of $L^{*k^{\prime}}-\{p\}$, then prior to deleting $q$, $p$ was above layer k1k-1, which is a contradiction.